OpenStudy (anonymous):

Slope = -2, passing through (3,-1) I'm supposed to write the equation of the line that satisfies the given conditions. And I must write the answer in slope intercept form. I've gotten down to to -2 = (-2)(4) + b, but I don't know how to solve for b.

OpenStudy (anonymous):

y-(-1) = -2 * (x-3) y=-2x+5

OpenStudy (anonymous):

What does * mean?

OpenStudy (anonymous):

Inik gave you a simple numerical answer, however, if you want to know how to solve for b you simply take your slope intercept equation: y=mx+b Insert slope: y=(-2)x+b Insert the X and Y coordinate given (3,-1) -1=(-2)(3)+b Solve for b. -1=(-6)+b 5=b

OpenStudy (anonymous):

* is multiplication :)

OpenStudy (anonymous):

Oh ok. Thank you Inik, Jesse and Melor.